OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = 2x + 2 and passes through the point (5, 3)

OpenStudy (anonymous):

Y-y1=m(x-x1)

OpenStudy (anonymous):

so..........

OpenStudy (anonymous):

slope stays the same

OpenStudy (anonymous):

So you do y-3=2x-5 y-3=2x-10 (due to distributing) add 3 to both sides Answer Y=2x-7

OpenStudy (anonymous):

Understand or still need help??

OpenStudy (anonymous):

understood thanks!

OpenStudy (anonymous):

Your welcome!!